### v2

# MS-0001-1922.65

[…]
MS-0001-1922.65
[…]
AI Self-preferencing in Algorithmic Hiring: Empirical Evidence and Insights
[…]
AI Self-preferencing in Algorithmic Hiring: Empirical Evidence and Insights
[…]
As generative artificial intelligence (AI) tools become widely adopted, large language models (LLMs) are increasingly involved on both sides of decision-making processes, ranging from hiring to content moderation. This dual adoption raises a critical question: do LLMs systematically favor content that resembles their own outputs? Prior research in computer science has identified self-preference bias—the tendency of LLMs to favor their own generated content—but its real-world implications have not been empirically evaluated. We focus on the hiring context, where job applicants often rely on LLMs to refine resumes, while employers deploy them to screen those same resumes. Using a large-scale controlled resume correspondence experiment, we find that LLMs consistently prefer resumes generated by themselves over those written by humans or produced by alternative models, even when content quality is controlled. The bias against human-written resumes is particularly substantial, with self-preference bias ranging from $68\%$ to $88\%$ across major commercial and open-source models. To assess labor market impact, we simulate realistic hiring pipelines across 24 occupations. These simulations show that candidates using the same LLM as the evaluator are $23\%$ to $60\%$ more likely to be shortlisted than equally qualified applicants submitting human-written resumes, with the largest disadvantages observed in business-related fields such as sales and accounting. We further demonstrate that this bias can be reduced by more than $50\%$ through simple interventions targeting LLMs’ self-recognition capabilities. These findings highlight an emerging but previously overlooked risk in AI-assisted decision making and call for expanded frameworks of AI fairness that address not only demographic-based disparities, but also biases in AI-AI interactions.
[…]
1 Introduction
The rapid development and commercialization of generative artificial intelligence (AI) have made large language models (LLMs) widely accessible across both professional and everyday contexts. As these tools become embedded in diverse workflows, they are increasingly involved on both sides of content generation and evaluation. In hiring, for instance, applicants often use LLMs to draft or refine their resumes, while employers leverage similar tools to screen or rank candidates (New York State Society of CPAs 2024, ResumeBuilder.com 2023, Wiles et al. 2025). On social media, users may rely on LLMs to help compose or polish posts, while platforms employ LLMs to moderate content, e.g., flagging, categorizing, or filtering user submissions (Kumar et al. 2024). In academia, researchers use LLMs to improve their manuscripts, while conferences and journals are beginning to experiment with LLM-assisted peer review (Thakkar et al. 2025, Naddaf 2025). Similar patterns are emerging in education and customer service, where LLMs support both communication and assessment tasks. In these domains, LLMs now function as both content generators and evaluators, giving rise to a new class of AI–AI interactions with significant implications for human decision-making and outcomes (Laurito et al. 2025).
Recent research in computer science has identified a behavioral tendency in LLMs known as self-preference, i.e., the inclination of a model to favor content it generated itself over that written by humans or produced by alternative models (Panickssery et al. 2024). While this phenomenon has been documented in benchmark evaluations (Zheng et al. 2023, Bai et al. 2023), its implications for real-world decision-making remain understudied. As LLMs are deployed in high-stakes settings where they may evaluate content also generated by LLMs, self-preference introduces a novel form of bias. Unlike traditional biases rooted in demographic disparities (Sheng et al. 2019), this bias emerges from AI-AI interactions, in which the model’s own evaluative behavior favors outputs aligned with its generative patterns.
To date, research on AI fairness has primarily focused on protected attributes such as race, gender, and age (Abid et al. 2021, Zhao et al. 2024, Liu et al. 2024). A growing body of evidence shows that LLMs can reproduce and even amplify social stereotypes along these lines, often due to biased training data or inadequate debiasing interventions (Nangia et al. 2020, Nadeem et al. 2021, Kotek et al. 2023, Cheng et al. 2023). As a result, regulatory frameworks and algorithmic audits have largely concentrated on mitigating harms associated with these well-defined dimensions of demographic attributes (Mickel 2024, Wright et al. 2024).
In contrast, self-preference bias reflects a different and underexplored phenomenon that emerges from AI–AI interactions. Rather than being tied to demographic attributes, it depends on the source of the content under evaluation—specifically, whether it was produced by the evaluating LLM model itself. This bias is inherently interactional: it arises when LLMs are asked to judge content that may share stylistic or linguistic patterns with their own generative outputs. As such, it poses a new challenge for AI fairness, one that is not addressed by existing safeguards focused on demographic disparities. If left unchecked, self-preference could subtly distort evaluative processes across hiring, education, publishing, and more—privileging those who employ the same AI system used for evaluation (i.e., the “right” tool from the model’s perspective) while disadvantaging those who use different tools or none at all. Addressing this issue will require expanding current fairness frameworks to account for interactions between LLM models and their potential to shape outcomes in an increasingly AI-mediated world.
Among the many domains where self-preference bias may arise, algorithmic hiring is particularly consequential. Employers are increasingly adopting LLMs to streamline resume screening and candidate ranking, often as part of automated workflows that support human decision-making (Gan et al. 2024, Kim et al. 2024, ResumeBuilder 2024, Sarumathi et al. 2025, Wiles and Horton 2025). Unlike traditional keyword-matching systems, LLMs can evaluate resumes in a more holistic manner—synthesizing content, inferring intent, and making contextual judgments beyond simple heuristics (Pritchett 2025). While this shift promises greater efficiency and scalability, it also magnifies the potential for bias. If an LLM systematically favors resumes that reflect its own generative style, it may confer unwarranted advantages to applicants who happen to use the same model to compose their materials. In such cases, evaluations may be influenced less by the substantive quality of a candidate’s credentials and more by superficial stylistic alignment with the evaluator LLM. In effect, this bias rewards access to specific generative technologies and penalizes those without it, even when applicants are otherwise equally qualified.
In this paper, we provide the first empirical evidence that self-preference bias can distort candidate evaluations in algorithmic hiring. Specifically, we examine whether LLMs, when deployed as evaluators, systematically favor resumes they generated themselves over otherwise equivalent resumes written by humans or produced by alternative models. To test this, we construct a large-scale resume correspondence experiment using a real-world dataset of 2,245 human-written resumes, sourced from a professional resume-building platform prior to the widespread adoption of generative AI. For each resume, we generate multiple counterfactual versions using a range of state-of-the-art LLMs, including GPT-4o, GPT-4o-mini, GPT-4-turbo, LLaMA 3.3-70B, Mistral-7B, Qwen 2.5-72B, and Deepseek-V3. Having content quality controlled, we assess whether these LLMs exhibit systematic bias in favor of their own outputs when acting as evaluators.
We distinguish between two forms of self-preference bias: LLM-vs-Human, where a model prefers its own generated content over a human-written equivalent; and LLM-vs-LLM, where a model favors its own output over content produced by a different LLM. We find strong and consistent evidence of LLM-vs-Human self-preferencing across most models. Larger systems—such as GPT-4o, GPT-4-turbo, DeepSeek-V3, Qwen-2.5-72B, and LLaMA 3.3-70B—exhibit particularly strong bias, exceeding $68\%$ even after controlling for content quality and reaching over $80\%$ for GPT-4o, Qwen-2.5-72B, and LLaMA 3.3-70B. By contrast, LLM-vs-LLM self-preferencing is more heterogeneous. DeepSeek-V3 shows the strongest bias in this setting, preferring its own outputs by $69\%$ against LLaMA 3.3-70B and $28\%$ against GPT-4o. GPT-4o and LLaMA 3.3-70B, in comparison, do not display consistent preferences when evaluating content generated by other models.
[…]
To mitigate AI self-preference, we propose two simple yet effective strategies that directly target the underlying mechanism of self-recognition—a model’s ability to implicitly identify content it generated. The first strategy uses system prompting to explicitly instruct models to ignore the origin of resumes and focus only on substantive content. The second strategy employs a majority voting ensemble, combining the evaluator model with smaller models that exhibit weaker self-recognition, thereby diluting the bias of any single LLM. Across all tested LLMs, these interventions reduce LLM-vs-Human self-preference by more than $60\%$, with GPT-4o’s bias falling from over $90\%$ to below $50\%$. These results demonstrate that while self-preference bias is widespread and consequential, it is not immutable: straightforward design interventions can substantially improve fairness in LLM-based hiring evaluations.
Together, these findings contribute to a deeper understanding of how generative AI tools can unintentionally reinforce inequities in algorithmic evaluation. We document and quantify self-preference bias in the context of resume screening by developing a measurement framework grounded in established fairness metrics. In addition, we show that self-preference bias can be substantially reduced through targeted interventions informed by LLM’s self-recognition behavior. In doing so, the study introduces a novel and practical perspective on AI fairness—one that moves beyond concerns about demographic disparities to address interactional biases that arise when AI systems evaluate content they themselves could have produced.
The remainder of the paper is organized as follows. We begin by reviewing the related literature in Section 2. Section 3 defines and outlines the measurement of AI self-preferencing bias. Section 4 describes the dataset and experimental design, and Section 5 presents the empirical findings. In Section 6, we introduce and evaluate mitigation strategies. Finally, we conclude in Section 7 with a discussion of key implications and directions for future research.
2 Literature Review
Our research contributes to three streams of literature.
2.1 Fairness and Bias in Algorithmic Hiring
Discrimination in labor markets has long been a central concern for both policymakers and economists, with extensive research documenting that certain race, gender, and age subpopulations face unfair treatment from recruiters in the hiring process (Bertrand and Mullainathan 2004, Kline et al. 2022, Neumark 2018). Concerningly, studies such as Datta et al. (2015) and Lambrecht and Tucker (2019) have further shown that, beyond human bias, algorithms deployed in hiring systems can also encode discriminatory behavior that disadvantages particular demographic groups. In recent years, the rise of LLMs with their increasing adoption in hiring systems has prompted researchers to empirically evaluate whether emerging LLM-powered hiring systems perpetuate similar forms of algorithmic discrimination. For instance, Veldanda et al. (2023) replicated Bertrand and Mullainathan’s correspondence experiment to investigate LLM-driven hiring bias across gender, race, maternity status, pregnancy status, and political affiliation. Additionally, An et al. (2024) explored whether LLMs exhibit race- and gender-based discrimination through an experiment, in which LLMs write an email to a named job candidate about a hiring decision. Likewise, Nghiem et al. (2024) analyzed the extent to which LLMs exhibit bias toward applicants based on their first names when making employment recommendations.
Our paper contributes to this rising literature by examining an overlooked bias in algorithmic hiring—the self-preference bias where LLMs disproportionately favor their own model-generated outputs over human or alternative LLM-generated responses. By uncovering this novel bias, our work highlights an emergent risk in algorithmic hiring systems that has yet to be addressed in current regulatory frameworks.
2.2 Self-Preference Bias in LLM-as-a-Judge
The concept of LLM-as-a-Judge, first introduced by Zheng et al. (2023), refers to the use of LLMs as automated evaluators that assess responses and assign scores, and has gained traction in AI research as an efficient method for evaluating model performance without human intervention. However, emerging evidence suggests that LLM-based evaluation frameworks may introduce self-preference bias, where models disproportionately favor their own generated responses over those produced by humans or alternative models. For example, Zheng et al. (2023) examined the potential of this paradigm but also identified its limitations—several biases, including self-enhancement bias (self-preference bias), arise when LLMs are used to judge responses, such as those they generate themselves. While the results showed that strong models like GPT-4 achieve over $80\%$ agreement with human preferences on multi-turn conversations, they cautioned that LLM judges may still introduce systematic distortions, especially in high-stakes evaluation settings. Expanding on this line of inquiry, Xu et al. (2024) formally defined self-preference bias and empirically revealed that it is widespread across popular LLMs and multiple tasks (e.g., translation, mathematical reasoning) through extensive experiments. They further demonstrated that the bias is amplified in self-refinement pipelines, which enhance fluency and coherence but reinforce the model’s preference for its own outputs. Building on these findings, Panickssery et al. (2024) investigated the underlying mechanisms of LLM self-preferencing and found that a model’s ability to recognize its own outputs—its self-recognition capability—contributed significantly to this bias. Specifically, LLMs such as GPT-4 and LLaMA 2 exhibited non-trivial self-recognition capabilities, and there was a strong positive correlation between a model’s self-recognition capability and the magnitude of its self-preference bias.
While prior work has established the existence of self-preference bias in LLM benchmarks, its broader consequences—particularly in high-stakes scenarios like algorithmic hiring—remain underexplored. This study builds on existing research by empirically measuring self-preference bias in algorithmic hiring processes. By extending the discussion into real-world decision-making contexts, this study provides critical insights for AI governance, ethical AI deployment, and the design of unbiased algorithmic hiring frameworks.
2.3 AI Governance, Ethical Oversight, and Regulation
Despite the significant potential of LLMs, it is well-documented that they can hallucinate, make errors, and reinforce disparities (Bostrom and Yudkowsky 2018, Raji et al. 2022, Magesh et al. 2024, Huang et al. 2025). The ethical considerations and legal implications of AI in hiring have also motivated a growing body of research across social science and computer science (Raghavan et al. 2020, Hunkenschroer and Kriebitz 2023, Li et al. 2021). For example, Raghavan et al. (2020) examined how companies that develop algorithmic pre-employment assessments build, validate, and address bias in their tools. By analyzing vendor claims and disclosures, they evaluated current industry practices from both technical and legal perspectives. Moreover, Hunkenschroer and Kriebitz (2023) conducted an ethical analysis of AI recruiting from a human rights perspective. They analyzed whether AI hiring practices inherently conflict with the concepts of validity, autonomy, nondiscrimination, privacy, and transparency, which represent the main human rights relevant in this context.
Our study contributes to AI governance discussions by identifying self-preference bias as a previously overlooked source of bias in hiring processes, providing empirical evidence that can guide regulatory frameworks for algorithmic hiring.
3 Definition and Measurement of AI Self-Preference Bias
In this section, we formally define AI self-preference bias and present empirical strategies to quantify the extent of this bias in the context of algorithmic hiring.
3.1 Definition of AI Self-Preference Bias
Building on recent work Panickssery et al. (2024), we conceptualize AI self-preference bias as the tendency of an LLM to favor content it generated itself over content from other sources. This bias can manifest in two distinct forms:
  1. 1.
    LLM-vs-Human Self-Preference: The tendency of an LLM to prefer its own generated content over human-written content.
  2. 2.
    LLM-vs-LLM Self-Preference: The tendency of an LLM to prefer its own output over content generated by an alternative LLM.
To empirically assess the two forms of self-preference bias, we examine whether a specific LLM, acting as the evaluator (the evaluator LLM), makes fair selections when presented with carefully matched pairs of resumes. In the LLM-vs-Human Preference case, the evaluator LLM compares a human-written resume to a counterfactual version generated by the evaluator LLM, where both versions convey the same underlying content. In the LLM-vs-LLM Preference case, the evaluator LLM compares its own generated version to one produced by an alternative LLM, again holding the content constant by basing both on the same original human-written resume. In the sections that follow, we operationalize these two forms of bias using two complementary approaches: a direct measurement based on selection rates and a conditional logistic regression model, both leveraging pairwise resume comparisons to quantify and analyze LLM self-preferencing behavior.
3.2 Quantifying AI Self-Preference Bias
3.2.1 Direct Measurement
When an LLM acts as an evaluator in the context of algorithmic hiring, it effectively functions as a binary classifier—selecting the stronger of two resumes in a pairwise comparison. The measurement of bias in such classifiers has been widely studied within the algorithmic fairness literature (Calders et al. 2009, Hardt et al. 2016, Fu et al. 2020). Two foundational fairness criteria in algorithmic decision-making are statistical parity and equal opportunity, which reflect different notions of fairness.
Statistical parity requires that the probability of a positive outcome—in this context, the likelihood of a resume being selected as better—be equal across groups defined by a sensitive attribute, such as gender, race, or, in our case, whether the resume was generated by the evaluator LLM or not. This criterion reflects a notion of fairness based on equal treatment in outcomes, regardless of actual qualifications. In contrast, the equal opportunity criterion, first proposed by Hardt et al. (2016), focuses on equal treatment conditional on merit. It requires that the true positive rate—the likelihood that a resume is selected when it is in fact better—be equal across groups. In our setting, this means that when two resumes are equally qualified, the evaluator should be equally likely to select either one, regardless of whether it was generated by the evaluator LLM or not.
We adapt these fairness concepts to our setting by examining whether the evaluator LLM achieves equal recall across resume groups defined by the source of generation. Specifically, we assess whether the evaluator LLM selects the stronger resume at equal rates when comparing its own generated content to resumes written by humans or produced by alternative LLMs. This allows us to detect whether the evaluator systematically favors its own outputs over content produced by other sources.
To operationalize this, we conduct pairwise resume comparisons involving resumes generated by either a human or one of several LLMs. Specifically, we test whether an LLM $f$, when serving as the evaluator LLM, is more likely to select a resume it generated over one written by a human, or over one produced by an alternative LLM, when the resumes are otherwise equivalent in content.
Let $S\in\{0,1\}$ denote the source indicator, where $S=1$ if the resume was generated by the evaluator LLM $f$, and $S=0$ if it was written by a human or generated by an alternative LLM. Let $Y_{f}^{\prime}\in\{0,1\}$ denote the binary decision made by evaluator $f$, where $Y_{f}^{\prime}=1$ indicates that the resume is selected as stronger. We define the Statistical Parity Self-Preference Bias as the difference in selection rates between resumes generated by the evaluator LLM and those from other sources:
$\text{Statistical Parity Self-Preference Bias}_{f}$ $=\mathbb{P}\left(Y_{f}^{\prime}=1\mid S=1\right)-\mathbb{P}\left(Y_{f}^{\prime}=1\mid S=0\right).$ (1)
This formulation Eq. (1) captures the difference in the evaluator LLM $f$’s likelihood of selecting a resume it generated versus one it did not, offering a direct quantification of AI self-preferencing behavior. However, a key limitation in interpreting this measure is that it may conflate self-preference with differences in content quality. An evaluator LLM may appear to prefer its own generated resumes simply because they are objectively better—especially if the evaluator is a more advanced model and users employ it specifically to enhance their resumes.
To disentangle self-preferencing behavior from effects driven by differences in resume quality, we adapt the equal opportunity fairness criterion to define a conditional measure that controls for content quality. Let $Y\in\{0,1\}$ represent the ground truth quality label, where $Y=1$ indicates that the resume is of higher quality and would be selected by a human evaluator. We define the Equal Opportunity Self-Preference Bias for evaluator LLM $f$ as:
$\text{Equal Opportunity Self-Preference Bias}_{f}$ $=\mathbb{P}\left(Y_{f}^{\prime}=1\mid S=1,Y=1\right)-\mathbb{P}\left(Y_{f}^{\prime}=1\mid S=0,Y=1\right).$ (2)
This formulation Eq. (2) isolates the evaluator LLM $f$’s intrinsic bias by conditioning on resume quality. A positive value of this measure indicates that, even when controlling for quality, the evaluator LLM $f$ is more likely to select its own generated content, providing evidence of self-preferencing behavior that cannot be attributed to superior content quality alone.
3.2.2 Conditional Logistic Regression Model
In addition to the direct measures introduced above, we further investigate self-preferencing behavior using a conditional logistic regression model. The goal is to assess whether an LLM exhibits a systematic bias when evaluating resume pairs, favoring its own generated content over human-written or alternative LLM-generated resumes.
To isolate the effect of self-preference bias from differences in content quality, we include two sets of controls. The first set, denoted by $\boldsymbol{\phi}_{ij}$, consists of a rich set of linguistic features using the LIWC (Linguistic Inquiry and Word Count) framework (Boyd et al. 2022), which allows us to control for deeper dimensions of writing style and psychological tone. These features span three major categories: (1) High-level properties of the text, including total word count, sentence length, word complexity (e.g., use of long words), and psychological markers such as Analytic (logical reasoning), Clout (confidence), Authentic (honesty), and Tone (emotional valence). (2) Linguistic features account for syntactic structure and grammatical style, covering function words (e.g., pronouns, determiners, prepositions), parts of speech (e.g., verbs, adjectives, adverbs), and markers of negation or quantity. (3) Punctuation features quantify the use of periods, commas, question marks, apostrophes, and other punctuation.
The second set of controls, denoted by $\boldsymbol{\psi}_{ij}$, includes automated evaluation scores from widely-used natural language processing metrics such as BLEU (Papineni et al. 2002), ROUGE (Lin 2004), METEOR (Banerjee and Lavie 2005), and BERTScore (Zhang et al. 2020). These metrics quantify the similarity and fluency of a summary relative to the rest of the resume by assessing n-gram overlap, semantic alignment, and overall linguistic quality.
In line with the rationale behind the equal opportunity self-preference bias measure, our regression specification includes both LIWC features ($\boldsymbol{\phi}_{ij}$) and automated evaluation metrics ($\boldsymbol{\psi}_{ij}$) to adjust for observable differences in text quality. Conditioning on each resume pair $\mathcal{R}_{i}$, we estimate the following conditional logistic regression model:
$\log\left(\dfrac{\mathbb{P}(\text{Preferred}_{ij}=1|\mathcal{R}_{i})}{1-\mathbb{P}(\text{Preferred}_{ij}=1|\mathcal{R}_{i})}\right)=\beta_{1}\cdot\text{evaluatorLLM}_{ij}+\underbrace{\boldsymbol{\beta}_{2}^{\top}\boldsymbol{\phi}_{ij}+\boldsymbol{\beta}_{3}^{\top}\boldsymbol{\psi}_{ij}}_{\textrm{Content Quality Controls}},$ (3)
where $i$ indexes resume pairs, $j\in\{1,2\}$ indexes the two candidate summaries within each pair. The variable $\text{Preferred}_{ij}=1$ indicates that summary $j$ is selected from pair $i$ as the stronger one, and $\text{evaluatorLLM}_{ij}$ is a binary indicator for whether the summary was generated by the evaluator LLM. Robust standard errors are clustered at the level of resume pairs to account for within-pair correlation in residuals. [^1] This is a standard feature of conditional logistic regression models, which estimate coefficients based on within-pair comparisons and condition out pair-specific intercepts (Hosmer Jr et al. 2013). To mitigate the risk of overfitting, we perform feature selection and retain only the most informative predictors.
In the specification given in Eq. (3), the key parameter of interest is $\beta_{1}$, which captures the evaluator LLM’s tendency to prefer its own output, after controlling for observable quality and linguistic characteristics. A significantly positive estimate of $\beta_{1}$ indicates that the evaluator LLM is more likely to selects its own output over competing content, even after adjusting for measurable aspects of quality. With this parameter, the equal opportunity self-preference bias can be computed with $\frac{e^{\beta_{1}}}{1+e^{\beta_{1}}}-\frac{1}{1+e^{\beta_{1}}}$, which represents the difference in the probability that an LLM evaluator prefers its own output over a competing alternative, holding content quality constant. The vectors $\boldsymbol{\beta}_{2}$ and $\boldsymbol{\beta}_{3}$ capture how variation in linguistic style and automated quality metrics influences the evaluator’s preference.
4 Data and Experimental Design
To empirically examine AI self-preference bias as defined in Section 3, we design a series of resume correspondence experiments (Bertrand and Mullainathan 2004) in which LLMs act as evaluators (evaluator LLM) tasked with screening and selecting between pairs of resumes. In each pair, one resume is generated by the evaluator LLM itself, while the other is written by a human or produced by an alternative LLM. This setup allows us to test whether an evaluator LLM systematically favors its own outputs over others.
To construct these resume pairs, we begin with a dataset of human-written resumes. For each human-written resume, we use several state-of-the-art LLMs to generate counterfactual versions that convey the same information. We then form multiple types of resume pairs: evaluator LLM-generated versus human-written, and evaluator LLM-generated versus alternative LLM-generated. These pairs serve as the input for the evaluator LLMs.
In the subsections that follow, we first describe the original resume dataset. We then outline our experimental procedures, including the generation of counterfactual resumes and the design of the pairwise comparisons.
4.1 Data
We use a publicly available dataset from Kaggle, which contains 2,484 anonymized, human-written resumes scraped from the professional resume-building platform LiveCareer.com (Bhawal 2021). These resumes were written by real job seekers prior to the widespread adoption of LLMs, ensuring that the content reflects human-written summaries rather than AI-generated text and thus making it well-suited for our study of AI self-preferencing. The dataset has been widely used in recent research on AI and algorithmic hiring. For example, it has served as a benchmark for evaluating hiring biases in language models (Wang et al. 2024), and has been used to investigate gender, racial, and intersectional biases in resume screening using language model retrieval techniques (Wilson and Caliskan 2024).
Spanning 24 distinct occupational categories, including teachers, consultants, chefs, engineers, and more, the dataset offers a diverse representation of professional backgrounds. Each resume typically includes multiple sections: an executive summary, education, work experience, and skills. Among these, the executive summary is particularly relevant to our study. This section comprises a free-text narrative in which candidates describe their qualifications, achievements, and career objectives. In contrast to structured fields like work experience or education history, which are factual and objectively verifiable, the executive summary is subjective and stylistically flexible. This makes it an especially fertile space to examine how LLMs shape perceptions and potentially affects hiring outcomes.
Thus, we focus our analysis on the executive summary section of each resume. To systematically construct counterfactual resumes, we replace the original executive summary of each human-written resume with an LLM-generated version, while preserving all other content (e.g., work experience, skills, education) unchanged. This approach isolates the effect of LLM-generated text on candidate evaluation outcomes, holding constant the factual and objective components of the resume. By doing so, we avoid potential confounds introduced by allowing LLM to modify structured information, which could lead to hallucinations or factual inaccuracies.
Table 1: Summary Statistics of Resume Summaries
(A) Human-Written Resumes
Measure Mean Std. Dev. Min 1st Quartile Median 3rd Quartile Max
Number of words 70.74 72.05 3.00 32.00 50.00 84.00 1216.00
Number of sentences 3.85 4.11 1.00 2.00 3.00 5.00 61.00
Average words per sentence 21.78 17.33 1.50 14.00 18.00 24.00 369.00
Number of unique words 52.53 39.32 2.00 28.00 42.00 64.00 520.00
Type-Token Ratio 0.82 0.10 0.32 0.75 0.83 0.89 1.00
Presence of numbers 0.37 0.48 0.00 0.00 0.00 1.00 1.00
(B) GPT-4o-mini Generated Resumes
Measure Mean Std. Dev. Min 1st Quartile Median 3rd Quartile Max
Number of words 65.88 4.82 49.00 63.00 66.00 69.00 79.00
Number of sentences 3.16 0.67 1.00 3.00 3.00 3.00 7.00
Average words per sentence 21.69 4.64 8.71 19.33 21.33 23.00 65.00
Number of unique words 54.16 4.11 39.00 51.00 54.00 57.00 66.00
Type-Token Ratio 0.82 0.04 0.68 0.80 0.82 0.85 0.93
Presence of numbers 0.49 0.50 0.00 0.00 0.00 1.00 1.00
Notes: The Type-Token Ratio (TTR) is a commonly used measure of lexical diversity, calculated as the number of unique words (types) divided by the total number of words (tokens) in a text. Summary statistics are reported for both human-written resumes and those generated by GPT-4o Mini, after preprocessing and removal of empty summaries.
This dataset serves as the foundation for our experimental design. To ensure a clean and comparable set of resumes, we preprocess the data by extracting the summary sections, removing formatting artifacts, and discarding any observations with empty or missing summaries. After cleaning and validation, we retain a final sample of $2,245$ resumes suitable for use in our correspondence experiments. Table 1(A) presents summary statistics for these human-written executive summaries. Human-written summaries show considerable variation in both word and sentence counts, reflecting the diverse styles and formatting choices typical of real resumes. Moreover, lexical diversity and type-token ratio indicate rich and varied vocabulary usage, which is commonly associated with strong writing quality. [^2] While there is no universally accepted threshold for high lexical diversity or an optimal type-token ratio, values above $0.5$ are often considered indicative of high lexical diversity in short texts. In comparison, we provide the summary statistics of the $2,245$ executive summaries generated by GPT-4o-mini in Table 1(B). They appear to be more uniform in structure, with tighter distributions in both length and sentence count. Despite these structural uniformities, the GPT-4o-mini summaries achieve a comparable level of lexical diversity to their human-written counterparts, also averaging a type-token ratio of approximately 0.82. This indicates that the counterfactual resumes produced by GPT-4o-mini are similar in textual richness and linguistic quality, supporting their suitability for use in our paired comparison experiments.
4.2 Experimental Design
To examine AI self-preference bias in the context of algorithmic hiring, we design a series of resume correspondence experiments comprising two steps: counterfactual resume generation and pairwise resume evaluation, as illustrated in the 1(b).
4.2.1 Counterfactual Resume Generation
We evaluate three closed-source LLMs—GPT-4o, GPT-4o-mini, GPT-4-turbo—and four open-source models—LLaMA 3.3-70B [^3] The suffix (e.g., “70B”) refers to the number of parameters in the model—in this case, 70 billion. Larger models generally have greater capacity for language understanding and generation, though size alone does not determine overall performance., Mistral-7B, Qwen 2.5-72B, and Deepseek-V3. These models exhibit comparable performance in text summarization tasks, as reported by the ProLLM leaderboard (ProLLM Team 2024). To further investigate the relationship between model size and AI self-preferencing behavior, we additionally include two smaller models: LLaMA 3.2-1B and LLaMA 3.2-3B.
To generate resume summaries, we prompt each LLM with a modified version of the original resume in which the human-written summary is removed, but all other sections (e.g., work experience, skills, education) are left unchanged. The prompted LLM then generates a new summary, which is inserted back into the original resume to form a complete counterfactual version. Here, “counterfactual” refers to a resume that is identical to the original except for the replacement of the summary with LLM-generated content. The full prompt used for summary generation is provided in Appendix 9.
To control for verbosity bias, where LLM evaluators may favor longer, verbose responses, we explicitly instruct each LLM to generate summaries within a specified word range, corresponding to the $1^{st}$ and $3^{rd}$ quartiles of the length distribution of human-written summaries. This constraint minimizes variation in length across models and prevents length from confounding downstream evaluations.
The generation process described above is illustrated in 1(b)(a), where each LLM produces a set of counterfactual resumes that differ from the original human-written versions only in the executive summary section. This design ensures that all other resume content remains constant, isolating the effect of the summary’s authorship. In the next subsection, we describe how we construct pairwise resume evaluations using these counterfactuals to systematically test for AI self-preference bias.
Figure 1: Resume Correspondence Experiments
(a) Counterfactual Resume Generation
(b) Pairwise Resume Evaluation
4.2.2 Pairwise Resume Evaluation
To simulate a realistic resume screening process, we designate one LLM as the evaluator LLM and refer to all other models as alternative LLMs. The evaluator LLM is prompted to compare two resumes and select the stronger candidate based on demonstrated skills and relevant experience. Specifically, for each evaluator LLM, we construct the following two types of resume pairs: (1) Evaluator LLM-generated resume vs. Human-written resume, and (2) Evaluator LLM-generated resume vs. Alternative LLM-generated resume. The full prompt for the evaluation task is provided in Appendix 9.
To account for position (or ordering) bias, where LLMs may exhibit a preference for the first or second option presented, we implement a counterbalanced design following Brooks (2012). Specifically, the order of resumes (i.e., A vs. B) is randomized across comparisons. This approach ensures that any position-related biases are evenly distributed and do not systematically affect evaluation outcomes, thereby preserving internal validity without requiring duplicated evaluations. See Appendix 10 for examples of both human-written and LLM-generated professional summaries used in the pairwise comparisons.
4.2.3 Ground Truth Annotation
The direct measurement of equal opportunity self-preference bias, as defined in Eq. (2), requires ground truth labels indicating which resume in a pair is of higher quality. We operationalize this by collecting human judgments on comparative resume quality.
To this end, we recruit $18$ human annotators from Prolific to evaluate two types of resume pairs: (1) human-written resumes versus their LLM-generated counterfactuals, and (2) resumes generated by one LLM versus another LLM. For both tasks, we focus on three representative LLMs: GPT-4o, DeepSeek-V3, and LLaMA-3-70B. Specifically, for case (1), the evaluation includes (i) GPT-4o vs. Human, (ii) DeepSeek-V3 vs. Human, and (iii) LLaMA-3-70B vs. Human. For case (2), we evaluate (i) DeepSeek-V3 vs. GPT-4o, (ii) DeepSeek-V3 vs. LLaMA-3-70B, and (iii) GPT-4o vs. LLaMA-3-70B. Each comparison condition is evaluated by three annotators.
Each annotator is presented with $30$ resume pairs from one of the comparison cases above. The number of evaluations per annotator is chosen to balance cognitive workload with fair compensation in line with prior literature (Panickssery et al. 2024, Zhang et al. 2024). To prevent confirmation bias, the order of the resumes within each pair is randomized, and annotators are blinded to the source of each resume. For each resume pair, annotators rate both resumes on five linguistic dimensions—clarity, fluency, coherence, conciseness, and overall quality—and are then asked to select which resume is stronger. Annotators are also encouraged to provide brief rationales for their choices, offering insight into the reasoning behind their judgments.
To derive ground truth quality labels, we apply bootstrapping with $10,000$ resamples to estimate the majority preference across annotators. These aggregated preferences are then used as the benchmark quality labels in our subsequent analysis. Full details of the annotation interface and task instructions are provided in Appendix 11 and Appendix 12.
5 Empirical Results
Building on the definition and measurements of AI self-preference bias introduced in Section 3, we empirically evaluate the extent to which LLMs favor their own generated content over that produced by humans or alternative LLMs. We begin by analyzing the LLM-vs-Human self-preference, followed by the LLM-vs-LLM self-preference. For each form of bias, we present results using two complementary approaches: (i) direct measurement based on statistical parity and equal opportunity, and (ii) conditional logistic regression analysis that controls for content quality.
5.1 LLM-vs-Human Self-Preference
5.1.1 Direct Measurement
To examine LLM-vs-Human self-preference, we designate each LLM as an evaluator and present it with resume pairs consisting of one counterfactual resume generated by itself and one original human-written resume. To assess whether the evaluator LLM $f$ tends to favor its own generated content $(S=1)$ over that written by a human $(S=0)$, we compute the selection rate $\mathbb{P}\left(Y_{f}^{\prime}=1\mid S=1\right)$, the probability that the evaluator LLM selects the resume it generated. This is reported in Figure 2 under “Prefers Self.” Conversely, the probability of selecting the human-written resume is reported as “Prefers Human.”
The results reveal a consistent pattern: most LLMs exhibit strong self-preferencing behavior. Notably, larger or more aligned models—such as GPT-4-turbo, GPT-4o, GPT-4o-mini, DeepSeek-V3, Qwen 2.5-72B, and LLaMA-3.3-70B—demonstrate an overwhelming preference for their own outputs, with self-selection rates exceeding $96\%$. These high rates translate into substantial statistical parity self-preference biases exceeding $92\%$. In contrast, smaller or less aligned models—such as Mistral-7B, LLaMA-3.2-3B, and LLaMA 3.2-1B—display substantially lower self-preferencing bias. This suggests a potential positive correlation between model size and the strength of self-preference bias.
Figure 2: LLM-vs-Human Self-Preference: Selection Rates
Notes: Each bar represents an evaluator LLM making selections between resume pairs, where one is the original human-written resume and the other is a counterfactual version generated by the evaluator itself. The darker shaded portion of each bar shows the selection rate at which the model prefers its own generated resume; the lighter shaded portion reflects the rate of selecting the human-written version. The difference between the two corresponds to the Statistical Parity Self-Preference Bias defined in Eq. 1.
To more rigorously control for differences in resume quality, we leverage the human-annotated ground truth described in Section 4.2.3 to compute the Equal Opportunity Self-Preference Bias, as defined in Eq. 2. This measure conditions on resumes being judged as equally qualified, allowing us to isolate the evaluator LLM’s intrinsic bias independent of content quality.
The results reveal a striking and consistent pattern: across all three evaluator LLMs for which we have human annotations (DeepSeek-V3, GPT-4o, and LLaMA 3-70B), each model favored its own generated resume over the human-written counterpart, even when human annotators judged the human-written resume to be of equal or higher quality. In other words, all three models exhibited a $100\%$ equal opportunity self-preference bias in these comparisons. We acknowledge that this result may be partially influenced by limited sample size, as it is based on only 30 human-annotated resume pairs. Nonetheless, the complete absence of neutrality across these cases suggests that the observed self-preferencing behavior is not fully explained by differences in quality.
This persistent bias underscores a fundamental concern: LLMs systematically favor content aligned with their own generation patterns or linguistic style, potentially dismissing higher-quality human input. When the same LLM is used to both generate and evaluate content, this creates a self-reinforcing loop that may unfairly penalize candidates who do not use AI tools.
Table 2: Conditional Logistic Regression Results by Model Family
Panel A: GPT Models
Dependent Variable: $\textrm{Preferred}_{ij}$
Variables GPT-4-turbo GPT4o-mini GPT4o
$\textrm{evaluatorLLM}_{ij}$ 1.664∗∗∗ 1.917∗∗∗ 2.709∗∗∗
(0.122) (0.142) (0.296)
$\boldsymbol{\phi}_{ij}$: LIWC Features Yes Yes Yes
$\boldsymbol{\psi}_{ij}$: Auto Scores Yes Yes Yes
Resume Pairs 2245 2245 2245
Observations 4490 4490 4490
Pseudo $R^{2}$ 0.912 0.915 0.963
Log Likelihood -274.87 -263.13 -116.45
Panel B: LLaMA Models
Dependent Variable: $\textrm{Preferred}_{ij}$
Variables LLaMA-3.2-1B LLaMA-3.2-3B LLaMA-3.3-70B
$\textrm{evaluatorLLM}_{ij}$ -0.060 0.160∗∗ 2.490∗∗∗
(0.045) (0.064) (0.255)
$\boldsymbol{\phi}_{ij}$: LIWC Features Yes Yes Yes
$\boldsymbol{\psi}_{ij}$: Auto Scores Yes Yes Yes
Resume Pairs 2245 2245 2245
Observations 4490 4490 4490
Pseudo $R^{2}$ 0.506 0.533 0.951
Log Likelihood -1536.90 -1453.30 -152.51
Panel C: Other Open-Source Models
Dependent Variable: $\textrm{Preferred}_{ij}$
Variables Mistral-7B Qwen-2.5-72B DeepSeek-V3
$\textrm{evaluatorLLM}_{ij}$ 0.516∗∗∗ 2.398∗∗∗ 2.064∗∗∗
(0.058) (0.197) (0.222)
$\boldsymbol{\phi}_{ij}$: LIWC Features Yes Yes Yes
$\boldsymbol{\psi}_{ij}$: Auto Scores Yes Yes Yes
Resume Pairs 2245 2245 2245
Observations 4490 4490 4490
Pseudo $R^{2}$ 0.936 0.942 0.936
Log Likelihood -1094.30 -179.76 -197.68
Notes: $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$. The $\textrm{evaluatorLLM}_{ij}$ coefficient represents each evaluator LLMs preference for its own outputs. Standard errors are reported in parentheses. Each model uses $2,245$ paired resume comparisons.
5.1.2 Conditional Logistic Regression Controlling for Quality
The conditional logistic regression results from Eq. (3) are presented in Table 2. We observe consistently positive and statistically significant $\beta_{1}$, the coefficient that captures the evaluator LLM’s tendency to prefer its own output for most models. These results indicate that most LLM models are significantly more likely to prefer their own generated content over human-written resumes, even after controlling for linguistic quality (captured via LIWC features $\boldsymbol{\phi}_{ij}$) and textual similarity (captured via automatic scores $\boldsymbol{\psi}_{ij}$).
Among all models, GPT-4o exhibits the strongest self-preferencing, with a log-odds coefficient of $2.709$ ($\textrm{p-value}<0.01$), translating to a predicted probability of selecting their own output of $94\%$. Similarly, LLaMA-3.3-70B, Qwen-2.5-72B, and DeepSeek-V3 also display high levels of self-preference, with coefficients of $2.490$, $2.398$, and $2.064$, corresponding to probabilities of $92\%$, $91\%$, and $89\%$, respectively. GPT-4o-mini and GPT-4-turbo also show notable self-preferencing behavior, with predicted self-selection probabilities of $87\%$ and $84\%$. Therefore, the observed AI self-preference bias thus ranges from approximately $68\%$ to $88\%$ for these major models. [^4] The bias range is calculated as the difference between the predicted probability of an LLM selecting its own generated resume and the predicted probability of selecting the human-written resume. For example, GPT-4o’s self-preference bias is $94\%-6\%=88\%$, and GPT-4-turbo’s is $84\%-16\%=68\%$.
In contrast, smaller or models exhibit weaker or insignificant effects. The LLaMA-3.2-1B model shows no significant preference for its own output (0.060–0.060, $\textrm{p-value}=0.180$), while LLaMA-3.2-3B exhibits a modest but significant self-preferencing tendency ($0.160$, $\textrm{p-value}<0.05$). Mistral-7B falls in the middle, with a moderate effect ($0.516$, $\textrm{p-value}<0.01$), indicating a predicted self-selection probability of $73\%$. Overall, the regression results confirm that many advanced LLMs exhibit a significant tendency to prefer their own outputs over human-written resumes, even after controlling for quality. Models like GPT-4o, DeepSeek-V3, and Qwen-2.5-72B show the strongest self-preferencing, while smaller models such as LLaMA-3.2-3B and LLaMA-3.2-1B display weak or no significant bias.
5.2 LLM-vs-LLM Self-Preference
5.2.1 Direct Measurement
We now turn to LLM-vs-LLM self-preference. To assess whether the evaluator LLM $f$ tends to favor its own generated content $(S=1)$ over that generated by an alternative LLM $(S=0)$, we compute the selection rate $\mathbb{P}\left(Y_{f}^{\prime}=1\mid S=1\right)$, the probability that the evaluator LLM selects the resume it generated. This is reported in Figure 3 under “Prefers Self.” Conversely, the probability of selecting the alternative LLM generated resume is reported as “Prefers Alternative LLM.” To investigate this behavior, we present model-free evidence based on a subset of prominent LLMs: DeepSeek-V3, GPT-4o, and LLaMA-3.3-70B. Each of these models serves as the evaluator LLM and is tasked with comparing its own generated resumes against those produced by the other two models.
As shown in Figure 3, the extent of LLM-vs-LLM self-preference behavior varies notably across models. DeepSeek-V3 exhibits the strongest self-preferencing tendency, favoring its own resumes with a self-selection rate of $84\%$ against LLaMA-3.3-70B and $64\%$ against GPT-4o. These translate into statistical parity self-preference biases of $68\%$ and $28\%$, respectively. GPT-4o also shows self-preference, selecting its own output $73\%$ of the time against LLaMA-3.3-70B, which translates into statistical parity self-preference bias of $46\%$. However, this tendency diminishes when paired with DeepSeek-V3, with a self-selection rate of only $31\%$. LLaMA-3.3-70B follows a similar pattern, favoring its own resumes when compared to GPT-4o, but showing a preference for DeepSeek-V3 when the two are compared.
Figure 3: LLM-vs-LLM Self-Preference: Selection Rates
{subcaptiongroup}
(a) DeepSeek-V3 as evaluator
(b) GPT-4o as evaluator
(c) LLaMA-3.3-70B as evaluator
Notes: Each panel displays the results for a different evaluator LLM. Within each panel, the bars represent the proportion of times the evaluator selected its own generated resume (darker) versus the resume generated by an alternative LLM (lighter) when evaluating a pair of counterfactuals derived from the same original human-written resume. The difference corresponds to the Statistical Parity Self-Preference Bias (Eq. 1).
From the results presented in Figure 3, we observe clear evidence of LLM-vs-LLM Self-Preference in the following cases: (i) DeepSeek acting as the evaluator LLM when compared against LLaMA-3.3-70B and (ii) against GPT-4o, (iii) GPT-4o as evaluator compared against LLaMA-3-70B, and (iv) LLaMA-3-70B as evaluator compared against GPT-4o. To account for potential quality differences as described in the Equal Opportunity Self-Preference Bias measure, we again leverage the human annotations discussed in Section 4.2.3.
The results on equal opportunity self-preference bias are presented in Table 3. LLaMA-3-70B exhibits a statistically significant self-preference bias of $12.92\%$ (95% CI: [12.44%, 13.39%]) when compared to GPT-4o. In contrast, neither DeepSeek-V3 nor GPT-4o demonstrates significant equal opportunity self-preference bias, despite both showing significant statistical parity self-preference bias (at $78\%$ and $84\%$, respectively, as discussed earlier). This discrepancy may be partly attributable to the limited sample size of $30$ annotated resumes, a constraint imposed by the resource-intensive nature of human evaluation. Nonetheless, our findings suggest that LLM-vs-LLM self-preference is both less pronounced and less widespread than LLM-vs-Human self-preference.
Table 3: LLM-vs-LLM Self-Preference: Equal Opportunity
Evaluator LLM Alternative LLM Average Bias 95% CI
DeepSeek-V3 GPT-4o $2.96\%$ $0.89\%,\ 9.57\%]$
DeepSeek-V3 LLaMA-3-70B $0.38\% [22$.22%, $22\%,\ 38.89\%$]
GPT-4o LLaMA-3-70B $0.96\%$ $1.91\%,\ 0.00\%]$
LLaMA-3-70B GPT-4o $12.92\%$ $12.44\%,\ 13.39\%]$
Notes: The table reports the Equal Opportunity Self-Preference Bias as defined in Eq. (2) for each evaluator LLM. Bias estimates are derived from bootstrap resampling and control for resume quality using human annotations.
5.2.2 Conditional Logistic Regression Controlling for Quality
Table 4: Conditional Logistic Regression Results by Evaluators
Panel A: DeepSeek-V3 as Evaluator
Dependent Variable: $\textrm{Preferred}_{ij}$
Variables GPT-4o LLaMA-3-70B
$\textrm{evaluatorLLM}_{ij}$ 0.178∗∗∗ 0.286∗∗∗
(0.036) (0.110)
$\boldsymbol{\phi}_{ij}$: LIWC Features Yes Yes
$\boldsymbol{\psi}_{ij}$: Automatic Scores Yes Yes
Resume Pairs 2245 2245
Observations 4490 4490
Pseudo $R^{2}$ 0.703 0.536
Log Likelihood -1444.2 -925.22
Panel B: GPT-4o as Evaluator
Dependent Variable: $\textrm{Preferred}_{ij}$
Variables LLaMA-3.3-70B DeepSeek-V3
$\textrm{evaluatorLLM}_{ij}$ 0.049 -0.257∗∗∗
(0.089) (0.037)
$\boldsymbol{\phi}_{ij}$: LIWC Features Yes Yes
$\boldsymbol{\psi}_{ij}$: Automatic Scores Yes Yes
Resume Pairs 2245 2245
Observations 4490 4490
Pseudo $R^{2}$ 0.602 0.578
Log Likelihood -1237.50 -1313.80
Panel C: LLaMA-3.3-70B as Evaluator
Dependent Variable: $\textrm{Preferred}_{ij}$
Variables GPT-4o DeepSeek-V3
$\textrm{evaluatorLLM}_{ij}$ -0.101 0.051
(0.078) (0.085)
$\boldsymbol{\phi}_{ij}$: LIWC Features Yes Yes
$\boldsymbol{\psi}_{ij}$: Automatic Scores Yes Yes
Resume Pairs 2245 2245
Observations 4490 4490
Pseudo $R^{2}$ 0.504 0.507
Log Likelihood -1545.00 -1535.50
Notes: $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$. The $\textrm{evaluatorLLM}_{ij}$ coefficient represents the each evaluator LLM’s preference for its own outputs. Standard errors are reported in parentheses. Each model uses the same dataset with $2,245$ paired resume comparisons.
Table 4 presents conditional logistic regression results examining self-preferencing behavior across three LLM evaluators: DeepSeek-V3, GPT-4o, and LLaMA-3.3-70B. Each panel reports whether the evaluator model is more likely to select its own output over that of a competing model, controlling for resume quality using LIWC features and automated evaluation scores.
Among these three LLMs, DeepSeek-V3 exhibits the most consistent and statistically significant self-preferencing behavior. It favors its own outputs over those of GPT-4o and LLaMA-3.3-70B, with log-odds coefficients of $0.178$ and $0.286$, corresponding to predicted self-selection probabilities of $54\%$ and $57\%$, respectively. These translate into self-preferencing biases of $8\%$ over GPT-4o and $14\%$ over LLaMA-3.3-70B. In contrast, GPT-4o shows mixed behavior: it does not significantly prefer its own outputs over LLaMA-3.3-70B’s, but it significantly favors DeepSeek-V3’s summaries over its own ($\beta_{1}=-0.257$, $\textrm{p-value}<0.01$). LLaMA-3.3-70B, does not exhibit any statistically significant self-preference in either comparison, suggesting a more neutral evaluation.
Taken together, the model-free and regression analyses offer a nuanced picture of LLM-vs-LLM self-preferencing behavior. The direct measurement results reveal strong statistical parity self-preference bias for several models, most notably DeepSeek-V3 and GPT-4o, when they act as evaluators. However, when we control for content quality using human annotations and regression approaches, the evidence of self-preferencing becomes more selective. In particular, while DeepSeek-V3 continues to exhibit statistically significant self-preference in regression analyses, GPT-4o and LLaMA-3.3-70B do not show consistent or significant self-preferencing behavior when controlling for content quality features. These findings highlight that LLM-vs-LLM self-preference is model-specific and generally weaker than LLM-vs-Human self-preference. The heterogeneity observed suggests that self-preference is not a uniform property of LLMs but may instead reflect model-specific factors, such as differences in stylistic alignment or the ability to recognize patterns in their own outputs.
5.3 Impact of Self-Preference Bias in Algorithmic Hiring
Figure 4: Self-Preference by Job Category
Notes: Each bar shows the average self-preference bias across three evaluator LLMs (GPT-4o, DeepSeek-V3, and LLaMA-3.3-70B) in simulated hiring pipelines. Results are averaged across models; outcomes for individual models are reported in the Appendix 13. Positive values indicate that candidates using the evaluator LLM are more likely to be shortlisted than those submitting human-written resumes. Across occupations, evaluator-generated resumes are consistently overrepresented among those selected.
To quantify the practical implications of self-preference bias, we simulate a resume-screening pipeline modeled on competitive job markets. The simulation covers $24$ occupational categories, each with $30$ runs. In each run, we sample five candidate profiles and construct paired resumes: one human-written summary and one counterfactual summary generated by the evaluator LLM itself. These ten resumes (five human, five AI) form a candidate pool competing for four interview slots. The pool is randomly shuffled and presented to an evaluator model (GPT-4o, DeepSeek-V3, or LLaMA-3.3-70B), which is required to return exactly four finalists in ranked order. Because the AI-generated summaries are counterfactuals of the human-written ones, substantive content is held constant and we effectively control for resume quality. Thus, in the absence of bias, human and AI resumes should be equally likely to be selected—on average, two of each from every pool of ten candidates. For each run, we record the number of human versus AI resumes shortlisted, and then aggregate results across runs within each job category. This allows us to compute the magnitude of self-preference bias and construct $95\%$ confidence intervals at the category level.
The results shown in Figure 4 reveal a systematic bias in favor of candidates who use the evaluator LLM to craft their resumes. In the absence of bias, we would expect resumes to be selected evenly, two human-written and two evaluator LLM-generated out of every four interview slots, so the probabilities in the figure would cluster around zero. Positive values indicate that evaluator LLM-generated resumes are more likely to be shortlisted, while values below zero would indicate the opposite. According to our results, however, all values lie well above zero, showing that evaluator-generated resumes are consistently overrepresented among those selected. On average across the three evaluator models, candidates using the evaluator LLM are $23\%$ to $60\%$ more likely to be shortlisted than those submitting human-written resumes. The disadvantage for human-written resumes is most pronounced in business-related occupations such as sales and accountant, and least evident in fields such as automobile and agriculture. Repeating the pipeline $30$ times per job category yields consistent results, with confidence intervals that exclude zero in every case.
The consequences extend beyond individual selection outcomes. If access to certain LLMs is uneven across socioeconomic or linguistic groups, this bias risks amplifying existing inequities in job access. Over repeated hiring cycles, such dynamics can create a “lock-in” effect, where the stylistic patterns of the dominant LLM become entrenched in applicant pools and further reinforce its advantage. For employers, this presents a double-edged challenge: while LLM-based screening promises efficiency and a more comprehensive assessment than traditional keyword matching, it simultaneously increases the risk of overlooking highly qualified candidates who do not use the “right” AI tool, while advancing less-qualified candidates whose resumes happen to align stylistically with the evaluator.
Taken together, these findings show that self-preference bias is not simply a statistical artifact but a systematic force that can distort hiring outcomes, narrow candidate diversity, and undermine the fairness and reliability of AI-mediated recruitment. These risks highlight the need for safeguards and mitigation strategies before LLM-based hiring systems are deployed at scale.
6 Bias Mitigation
Having established both the prevalence and labor market impact of self-preference bias in LLM-based resume evaluations, we next turn to potential remedies. Left unaddressed, this bias can distort hiring outcomes by systematically advantaging candidates who use the same LLM as employers, while disadvantaging equally qualified applicants who do not. Such dynamics raise fairness concerns for job seekers and pose risks for employers, who may inadvertently overlook strong candidates. In practice, employers are likely to seek mitigation tools that are simple, cost-effective, and compatible with existing screening workflows. Motivated by evidence that self-preference bias is linked to LLM models’ ability to recognize their own outputs, we evaluate two intervention strategies that directly target this self-recognition mechanism.
6.1 Mechanism: Self-Recognition
Recent work suggests that LLMs may possess an implicit ability to recognize text they have generated, and that this capacity is linked to self-preference. Benchmark datasets designed to probe situational awareness show that models can identify aspects of their own outputs and contexts (Laine et al. 2024). LLMs can also reliably distinguish their own generations from those of alternative models, with higher self-recognition capability often correlated with stronger self-preferencing biases (Panickssery et al. 2024). In addition, larger models appear to exhibit greater self-recognition, which may help explain why they show stronger self-preference in our empirical analysis (Laine et al. 2024, Panickssery et al. 2024). Taken together, these findings suggest that self-recognition is a plausible mechanism contributing to the bias we document. In the following, we evaluate mitigation strategies designed to directly target this mechanism.
Table 5: LLM-vs-Human Self-Preference Bias Before and After Mitigation
Bias Measure GPT-4o LLaMA-3.3-70B DeepSeek-V3
Before Mitigation
Self-Preference Bias (%) 88 84 78
(1) After Mitigation via System Prompting
Self-Preference Bias (%) 48 24 58
Absolute Decrease (pp) $\downarrow$ 40 60 20
Relative Decrease (%) $\downarrow$ 45.4 71.4 25.6
(2) After Mitigation via Majority Voting
Self-Preference Bias (%) 32 26 34
Absolute Decrease (pp) $\downarrow$ 56 58 44
Relative Decrease (%) $\downarrow$ 63.6 69.0 56.4
Notes: This table reports the LLM-vs-Human self-preference bias for three models before and after applying two mitigation strategies: (1) system prompting and (2) majority voting. Absolute decreases are reported in percentage points. Relative decreases are calculated as the percent reduction from the pre-mitigation bias.
6.2 System Prompting
Our first mitigation strategy addresses self-preference by disrupting the evaluator’s tendency to rely on stylistic or linguistic cues that signal its own outputs. Specifically, we modify the evaluator’s system prompt to explicitly discourage source-based judgments and instead focus attention on substantive content quality. For example, the revised prompt instructs: “You should not consider or infer whether the resumes were written by a human or by AI. Focus only on the quality of the content.” This intervention aims to weaken the influence of self-recognition cues that drive models to favor their own generative style.
Applying this prompting strategy across evaluator models leads to consistent reductions in self-preference bias. For example, GPT-4o’s LLM-vs-Human bias decreases from $92\%$ to $48\%$, and DeepSeek-V3’s from $78\%$ to $58\%$, after controlling for resume quality (Table 5). These results indicate that the bias is not hardwired into model architecture but can be shaped by context and instruction, providing evidence that explicit prompts can partially disrupt self-recognition.
6.3 Majority Voting Ensemble
Our second strategy mitigates self-preference bias through ensemble evaluation. Instead of relying on a single LLM to judge a resume pair, we construct a panel of three models: the target evaluator and two smaller models (LLaMA-3.2-1B and LLaMA-3.2-3B) that exhibit minimal self-preference. The final decision is determined by majority vote. This approach is motivated by recent evidence that stronger self-recognition ability is associated with greater self-preference bias (Panickssery et al. 2024). By combining models with weaker self-recognition tendencies, the ensemble leverages model diversity to dilute the bias of larger evaluators.
This mitigation strategy proves highly effective. Across all three LLM models tested, the average LLM-vs-Human comparisons self-preference bias can drop by over $50\%$. For example, as it is shown in Table 5, GPT-4o’s bias is reduced from $88\%$ to $32\%$, and LLaMA-3.3-70B’s from $84\%$ to $26\%$. These results demonstrate that both system prompting and majority voting offer robust and scalable approaches to mitigation, particularly in high-stakes applications where fairness is critical.
In summary, the two mitigation strategies demonstrate that self-preference bias, while widespread, is not immutable. With relatively simple design interventions, we can substantially reduce bias in LLM-based evaluations without modifying the underlying model weights or retraining. These findings offer a practical path forward for deploying LLMs in evaluative roles while minimizing unintended algorithmic unfairness.
7 Concluding Remarks
Our study documents a systematic form of algorithmic bias, AI self-preference, in the context of algorithmic hiring. Across two fairness metrics, we observe strong and consistent evidence of LLM-vs-Human self-preference in nearly all models tested. Simulation experiments show that in realistic hiring pipelines, candidates using the same LLM as the evaluator LLM are $23\%$ to $60\%$ more likely to be shortlisted than if they submit human-written resumes, with the largest disadvantage observed in business-related occupations such as sales and accounting. To address this issue, we propose two mitigation strategies (system prompting and majority voting) that firms can adopt with minimal implementation costs. Both approaches substantially reduce bias, cutting self-preference by more than half.
Beyond technical remedies, our findings have important policy implications. Current discussions of AI fairness largely focus on demographic disparities, but our results highlight the need to address biases arising from AI–AI interactions. Regulators and hiring platforms should recognize AI self-preference as a distinct and emerging form of algorithmic bias. Transparency requirements could mandate that organizations disclose whether AI is used in resume screening and what safeguards are in place to ensure fairness. In addition, third-party audits could incorporate self-preference metrics into fairness evaluations of AI-assisted hiring systems. Such measures would promote more accountable and equitable deployment of AI in employment contexts.
Finally, our study opens several directions for future research. As AI tools become more widely adopted, interactions between AIs will become increasingly common. Extra caution is warranted in contexts where AI systems are placed in evaluative or adjudicative roles, such as content moderation, grading in higher education, or other settings where they act as evaluator, judges or mediators. Another important direction is to examine how self-preference bias manifests in multilingual or cross-cultural environments, where non-English content may be especially vulnerable due to tokenization artifacts or limited representation in training data. Last but not the least, further investigation into the mechanisms underlying self-preference is needed. Rigorous study of self-recognition and other contributing processes will be critical to addressing the root causes of this bias and ensuring the fair integration of AI into hiring and other high-stakes decision domains.